OpenStudy (anonymous):

A wheel makes 1 3/10revolutions per minute. If it rotates for 85minutes, how many revolutions does it make?

jhonyy9 (jhonyy9):

so can you rewriting 1 3/10 without denominator ,not fractionaly ?

OpenStudy (anonymous):

im confuse

jhonyy9 (jhonyy9):

so than 1 3/10 = (1*10 +3)/10 = can you contiue it ?

OpenStudy (anonymous):

i will try.

OpenStudy (anonymous):

1.3

jhonyy9 (jhonyy9):

so than you know that 1,3 rev will be maked per minute so than how many will be maked in 85 minutes ? how do you think ?

OpenStudy (anonymous):

im confuse with this part

jhonyy9 (jhonyy9):

us the rule of 3 simple ,so than you can writing 1,3 rev ------- 1 minute x rev --------85 minutes ------------------ cross multiplie and will get 1*x = 85*1,3 x=85*1,3 = hope so much that you will can calcule it sure easy right yes ?